If P(A) is the probability that it's raining in Amherst, which of the following would equal the probability that it's not raining in Amherst?

a. 1 / P(A)

b. P(A) / 1

c. 1 - Pr(A)

d. P(A) - P(~A)

Suppose that at this time of year, it rains at some point in the day on about a quarter of the days. Before getting any additional evidence, what are the prior odds that it's going to rain at some point in the day?

a. 1 : 3

b. 1 : 4

c. 4 : 1

d. 3 : 1
Answer #1

1. Ans: c. 1 - Pr(A)

If P(A) is the probability of an event, then 1 - P(A) is the probability of the complement of the event.

2. P(rain) = 1/4

P(no rain) = 1 - 1/4 = 3/4

Odds that its going to rain = P(rain) : P(no rain)

= 1/4 : 3/4

= 1 : 3

Ans: a. 1 : 3